What is the APR (annual) for a loan that charges a $12 fee (interest) to borrow $100 for a loan period of 10 days.

(a) 312%

(b) 438%

(c) 120%

(d) 304%
Answer #1

interest rate for 10 days = 12/100 = 12%

APR = number of periods * rate per period

= 365/10 * 12%

= 438%

choose b)